# The Fate of Entanglement

### v4

@@ II Separable set
where $\rho_{j}$ is a physical density matrix for subsystem $j$. The set of separable sates is convex [6], namely states composed of a mixture of product states, $\rho_{\rm sep}=\sum_{k}p_{k}\rho_{\rm prod}^{(k)}$ with $p_{k}\geqslant 0$ and $\sum_{k}p_{k}=1$, are also un-entangled. In the space of all physical states, separable ones thus form a “continent” surrounded by an “ocean” of entangled states, see Fig. 1b. We stress however that this representation does not account for the intricate geometry of the boundary, but it illustrates the relevant ingredients for our discussion, namely the fact that the set of separable states is convex and has a non-vanishing volume. In our schematic representation, highly-entangled states live in the deep-blue regions, whereas in the center of the separable continent lies the maximally-mixed state, i.e., the identity.
@@ II.1 Biseparable set and genuine multipartite entanglement
where $\mathcal{I}_{k}\cup\overline{\mathcal{I}}_{k}$ are bipartitions labelled by $k$, $p_{k}\geqslant 0$, and $\sum_{k}p_{k}=1$. As an example, for $m=3$ parties, the possible bipartitions are $(1,23)$, $(12,3)$ and $(13,2)$. Biseparable states contain at most $(m-1)$-party entanglement, which is not genuinely multipartite. Similarly to separable states, biseparable ones form a convex set which includes the separable continent. Using the same arguments as above, we conclude that during typical generalized evolutions $\rho(s)$, the state irreversibly loses GME at a stage prior to the sudden death of entanglement, implying that non-GME and bipartite entanglement are more robust than GME.
@@ II.3 Time
In a dynamical situation, the parameter $s$ is the time $t$, and we take the final state to be the state at some time $t_{f}$. In dynamical protocols where the system reaches a final state $\rho_{f}$ which lies inside the separable continent, one can conclude that there is a sudden-death time $t^{*}_{m}$ at which all $m$-party entanglement is lost until $t_{f}$. As the structure of the final state depends on the type of dynamical evolution under study, let us describe an important case in more detail: a quantum quench. One prepares a closed system to be in an eigenstate of a given Hamiltonian $H_{I}$, and at some time the Hamiltonian is abruptly changed to a different one, $H$, resulting in non-trivial time evolution. We then study the state of a subregion $A$ of the entire system: the $m$-party state $\rho(t)$ is obtained by partially tracing over the unobserved part, $B$. When $A$ is small compared to its complement $B$, the state is expected to effectively thermalize at large times [20, 21, 22]. As discussed above, temperature tends to destroy entanglement, which means that in numerous quench protocols the subregion final state $\rho(t_{f})$ will be separable. Although not all quenches will lead to a sudden death of multipartite entanglement, one can ask about the typicality of such a fate. For instance, if the initial condition has an energy expectation value close to the groundstate of $H$, instead of thermalisation, oscillatory behavior with persistent entanglement can occur. However, only a minority of eigenstates of a local Hamiltonian $H$ have an energy near the minimum (or maximum) of the spectrum: the bulk of the levels lies near the middle of the spectrum. As such, a typical initial condition will inject enough energy so that the evolution will scramble the quantum information encoded within $A$. In many cases one can be more precise: the state of $A$ at large times can is accurately approximated by an effective thermal Gibbs state, $\operatorname{Tr}_{B}(\exp(-H/T_{\rm eff}))/\mathcal{Z}$, When the effective temperature $T_{\rm eff}$ is not small (which is typical), an entanglement sudden death for the subregion follows. Below we shall rigorously verify the dynamical sudden death with numerous examples.
@@
To complement the discussion of late-time dynamics, we can argue in full generality that the multiparty entanglement dynamics after a quench from a pure product state typically follows a rise-and-fall behavior. For early times, we consider the evolution of the system with parameter $s=t^{-1}$. The “final” state of this evolution is the initial state of the quench, namely a pure product state with rank one, i.e., not full rank. As we previously discussed, such states lie on the boundary, or the shore, of the separable continent. Therefore, there is no entanglement sudden death in this evolution. Looking back at the quench protocol with time as the parameter, this implies that entanglement is generated at $t=0^{+}$ after the quench. As discussed above, in numerous situations, the state typically lands on the continent at late times, or in very shallow waters at a finite time, yielding an entanglement sudden death or strong suppression. These early- and late-time behaviors result in the rise-and-fall dynamics of entanglement, which have been observed for simple entanglement measures in various systems [23, 24, 25, 26, 27, 28, 29].
@@ II.4 Space
We now study the fate of entanglement as a function of the separation between the $m$ subsystems by scaling their separations by $\lambda$, which parametrizes the evolution. At large $\lambda$, in a local physical system the state will factorise into a product form, where the description of each subsystem becomes independent, since all correlations become suppressed at large separations. If the asymptotic product state is of full rank, there exists a critical scale beyond which all entanglement vanishes. As a central application, we consider a quantum many-body system from which we extract the state of a subregion whose $m$ parts, $A=A_{1}\cdots A_{m}$, become more separated as $\lambda$ grows, $\rho(\lambda)$. The state at infinite $\lambda$ satisfies the product form, Eq. (1), where $\rho_{j}$ is the reduced density matrix of subsystem $j$ obtained by tracing out the complementary degrees of freedom. Let us take the entire system to be in equilibrium at a temperature $T~{}\!\!\geqslant~{}\!\!0$, described by the Gibbs state $\rho_{AB}=\mathcal{Z}^{-1}\exp(-H/T)$. At large separation, the state of subsystem $A_{1}$ is given by $\rho_{1}=\operatorname{Tr}_{A_{2}...A_{m}B}\sum_{n}p_{n}|E_{n}\rangle\langle E_{n}|$, where we have traced over the environment $B$, and the $(m-1)$ subsystems $A_{j}$. The sum runs over all eigenstates of the Hamiltonian $H$. The thermal Boltzmann probabilities are $p_{n}=e^{-E_{n}/T}/\mathcal{Z}$. We now need to determine whether $\rho_{1}$ has full rank. Let us first examine the restriction of the eigenstates of the Hamiltonian to $A_{1}$, $\rho_{1}^{(n)}=\operatorname{Tr}_{A_{2}...A_{m}B}|E_{n}\rangle\langle E_{n}|$. From the point of view of $A_{1}$, the degrees of freedom in the complement act as a bath, thus introducing statistical randomness in $\rho_{1}^{(n)}$. Since we take $A_{1}$ to be sufficiently small compared to the complement and the interactions generic (not fine-tuned), the bath generically has sufficient resources to induce statistical fluctuations that span the entire Hilbert space of $A_{1}$. Such ergodicity implies that $\rho_{1}^{(n)}$ will be of full rank. For instance, it is readily seen to occur for the majority of eigenstates as they live near the middle of the energy spectrum, and the eigenstate thermalization hypothesis [30, 31, 32] then states that the reduced density matrix on $A_{1}$ is approximately thermal, with the temperature determined by the energy $E_{n}$. The approximately thermal density matrix, obtained by restricting the Hamiltonian to $A_{1}$, then has full rank. Our statement, which we call the full-rank hypothesis (FRH), is more general: all eigenstates of a generic local Hamiltonian have a full-rank reduced density matrix associated with a sufficiently small subregion. In practice, the subregion should not exceed half the system. To complement the physical argument given above, we shall give an example in a frustrated quantum magnet in the next section.
@@ IV Icosahedral molecule
$H=J\sum_{{\rm bonds}\,\langle i,j\rangle}\sigma_{i}^{x}\sigma_{j}^{x}-h\sum_{{\rm sites}\,i}\sigma_{i}^{z}$ (6)
@@ IV.1 Full-rank hypothesis and zero temperature phase diagram
As a first step, we checked that the FRH holds for subregions of adjacent spins with 1, 2 and 3 sites by obtaining the $2^{12}$ eigenstates via exact numerical diagonalization, see Appendix B. We then considered the fate of entanglement with varying magnetic field at zero temperature, so that the evolution is parameterized by $s=h$, see Fig. 2b. At $h\!=\!0$, the 2-spin reduced density matrix is a full-rank separable state, which suggests the existence of a sudden-death value $h_{2}^{*}>0$ of the field below which the state is separable, and hence $\mathcal{E}(h<h_{2}^{*})=0$. We find $h_{2}^{*}\approx 0.6$. In contrast, the 3-spin reduced density matrix is not separable at small $h$ and we have $\mathcal{D}>0$ as shown in Fig. 2b. Turning to the GME between the spins on a triangular face, we find that it becomes very small for $h<0.65$, on the order of $10^{-6}$. Using the GTC, we are able to certify that the state is biseparable for $h\leqslant 0.25$. Although the certification becomes more difficult for larger field values, we believe that the true sudden death occurs near $0.65$.
@@ IV.4 Separation
with periodic boundary conditions such that $N+1\equiv N$. For our example, we randomly pick in the range $[-1,1]$ the 3 exchange couplings $J_{i}$, and 3 components of the Zeemann magnetic field $h_{i}$: $\vec{J}=(-0.443,0.0938,0.915)$ and $\vec{h}=(0.930,-0.685,0.941)$. We work with $N=14$ sites. The groundstate is unique with energy gap $\Delta E=2.0007$. We have verified that the FRH holds for all $2^{14}$ eigenstates by computing the smallest eigenvalue of the reduced density matrix for a subregion made of 4 adjacent spins, as shown in Appendix B, which confirms the generic nature of $H_{\rm xyz}$.
@@ VII Outlook
These new insights regarding the structure of entanglement in many-body states will have strong impact for quantum simulation and computation. For instance, efficient algorithms used to simulate quantum matter should incorporate the fact that distant particles are not entangled in spin/boson systems, which strongly constrains the variational space of many-body wave functions. Our work also paves the way for numerous outstanding research avenues. For instance, it will be essential to investigate how multipartite entanglement evolves in a plethora of realistic model Hamiltonians, and to determine the criteria that govern its demise under evolution. We have also seen examples of how quantum matter out-of-equilibrium generates stronger multipartite entanglement than in equilibrium. This is the tip of iceberg: we expect that non-equilibrium dynamics can lead to rich entanglement structures that await to be discovered.

### v5

@@ II Separable set
where $\rho_{j}$ is a physical density matrix for subsystem $j$. The set of separable sates is convex [6], namely states composed of a mixture of product states, $\rho_{\rm sep}=\sum_{k}p_{k}\rho_{\rm prod}^{(k)}$ with $p_{k}\geqslant 0$ and $\sum_{k}p_{k}=1$, are also un-entangled. In the space of all physical states, separable ones thus form a “continent” surrounded by an “ocean” of entangled states, see Fig. 1b. We stress however that this representation does not account for the intricate geometry of the boundary, but it illustrates the relevant ingredients for our discussion, namely the fact that the set of separable states is convex and has a non-vanishing volume. In our schematic representation, highly-entangled states live in the deep-blue regions, whereas in the center of the separable continent lies the maximally-mixed state, i.e., the identity. In the following, we focus on finite and discrete subsystems, and thus on finite-dimensional local density matrices. In the thermodynamic limit, the order of limits is chosen such that the total system remains incommensurably larger than the finite subsystems, which can in turn be large.
@@ II.1 Biseparable set and genuine multipartite entanglement
where $\mathcal{I}_{k}\cup\overline{\mathcal{I}}_{k}$ are bipartitions labelled by $k$, $p_{k}\geqslant 0$, and $\sum_{k}p_{k}=1$. As an example, for $m=3$ parties, the possible bipartitions are $(1,23)$, $(12,3)$ and $(13,2)$. Biseparable states contain at most (m1)(m-1)-party entanglement, which is not genuinely multipartite. Similarly to separable states, biseparable ones form a convex set which includes the separable continent. Using the same arguments as above, we conclude that during typical generalized evolutions $\rho(s)$, the state irreversibly loses GME at a stage prior to the sudden death of entanglement, implying that non-GME and bipartite entanglement are more robust than GME.
@@ II.3 Time
In a dynamical situation, the parameter $s$ is the time $t$, and we take the final state to be the state at some time $t_{f}$. In dynamical protocols where the system reaches a final state $\rho_{f}$ which lies inside the separable continent, one can conclude that there is a sudden-death time $t^{*}_{m}$ at which all $m$-party entanglement is lost until $t_{f}$. As the structure of the final state depends on the type of dynamical evolution under study, let us describe an important case in more detail: a global quantum quench. One prepares a closed system to be in an eigenstate of a given Hamiltonian $H_{I}$, and at some time the Hamiltonian is abruptly changed to a different one, $H$, resulting in non-trivial time evolution. We then study the state of a subregion $A$ of the entire system: the $m$-party state $\rho(t)$ is obtained by partially tracing over the unobserved part, $B$. When $A$ is small compared to its complement $B$, the state is expected to effectively thermalize at large times [20, 21, 22]. As discussed above, temperature tends to destroy entanglement, which means that in numerous quench protocols the subregion final state $\rho(t_{f})$ will be separable. Although not all quenches will lead to a sudden death of multipartite entanglement, one can ask about the typicality of such a fate. For instance, if the initial condition has an energy expectation value close to the groundstate of $H$, instead of thermalisation, oscillatory behavior with persistent entanglement can occur. However, only a minority of eigenstates of a local Hamiltonian $H$ have an energy near the minimum (or maximum) of the spectrum: the bulk of the levels lies near the middle of the spectrum. As such, a typical initial condition will inject enough energy so that the evolution will scramble the quantum information encoded within $A$. In many cases one can be more precise: the state of $A$ at large times is accurately approximated by an effective thermal Gibbs state, $\operatorname{Tr}_{B}(\exp(-H/T_{\rm eff}))/\mathcal{Z}$. When the effective temperature $T_{\rm eff}$ is not small (which is typical), an entanglement sudden death for the subregion follows. Below we shall rigorously verify the dynamical sudden death with numerous examples. Another important class of dynamical protocol are the local quenches, where the Hamiltonian is modified locally, for instance by connecting two different disjoint groundstates at $t=0$. In this case, the difference between the energy expectation value and the post-quench groundstate energy is not extensive, thus corresponding to a lower effective temperature in the steady state. It is an important question to understand whether this feature allows local quenches to generically escape the dynamical sudden death of entanglement, which we leave for future research.
@@
To complement the discussion of late-time dynamics, we can argue in full generality that the multiparty entanglement dynamics after a global quench from a pure product state typically follows a rise-and-fall behavior. For early times, we consider the evolution of the system with parameter $s=t^{-1}$. The “final” state of this evolution is the initial state of the quench, namely a pure product state with rank one, i.e., not full rank. As we previously discussed, such states lie on the boundary, or the shore, of the separable continent. Therefore, there is no entanglement sudden death in this evolution. Looking back at the quench protocol with time as the parameter, this implies that entanglement is generated at $t=0^{+}$ after the quench. As discussed above, in numerous situations, the state typically lands on the continent at late times, or in very shallow waters at a finite time, yielding an entanglement sudden death or strong suppression. These early- and late-time behaviors result in the rise-and-fall dynamics of entanglement, which have been observed for simple entanglement measures in various systems [23, 24, 25, 26, 27, 28, 29].
@@ II.4 Space
We now study the fate of entanglement as a function of the separation between the $m$ subsystems by scaling their separations by $\lambda$, which parametrizes the evolution. At large $\lambda$, in a local physical system the state will factorise into a product form, where the description of each subsystem becomes independent, since all correlations become suppressed at large separations. If the asymptotic product state is of full rank, there exists a critical scale beyond which all entanglement vanishes. As a central application, we consider a quantum many-body system from which we extract the state of a subregion whose $m$ parts, $A=A_{1}\cdots A_{m}$, become more separated as $\lambda$ grows, $\rho(\lambda)$. The state at infinite $\lambda$ satisfies the product form, Eq. (1), where $\rho_{j}$ is the reduced density matrix of subsystem $j$ obtained by tracing out the complementary degrees of freedom. Let us take the entire system to be in equilibrium at a temperature $T~\!\!\geqslant~\!\!0$, described by the Gibbs state $\rho_{AB}=\mathcal{Z}^{-1}\exp(-H/T)$. At large separation, the state of subsystem $A_{1}$ is given by $\rho_{1}=\operatorname{Tr}_{A_{2}...A_{m}B}\sum_{n}p_{n}|E_{n}\rangle\langle E_{n}|$, where we have traced over the environment $B$, and the (m1)(m-1) subsystems $A_{j}$. The sum runs over all eigenstates of the Hamiltonian $H$. The thermal Boltzmann probabilities are $p_{n}=e^{-E_{n}/T}/\mathcal{Z}$. We now need to determine whether $\rho_{1}$ has full rank. Let us first examine the restriction of the eigenstates of the Hamiltonian to $A_{1}$, $\rho_{1}^{(n)}=\operatorname{Tr}_{A_{2}...A_{m}B}|E_{n}\rangle\langle E_{n}|$. From the point of view of $A_{1}$, the degrees of freedom in the complement act as a bath, thus introducing statistical randomness in $\rho_{1}^{(n)}$. Since we take $A_{1}$ to be sufficiently small compared to the complement and the interactions generic (not fine-tuned), the bath generically has sufficient resources to induce statistical fluctuations that span the entire Hilbert space of $A_{1}$. Such ergodicity implies that $\rho_{1}^{(n)}$ will be of full rank. For instance, it is readily seen to occur for the majority of eigenstates as they live near the middle of the energy spectrum, and the eigenstate thermalization hypothesis [30, 31, 32] then states that the reduced density matrix on $A_{1}$ is approximately thermal, with the temperature determined by the energy $E_{n}$. The approximately thermal density matrix, obtained by restricting the Hamiltonian to $A_{1}$, then has full rank. Our statement, which we call the full-rank hypothesis (FRH), is more general: all eigenstates of a generic local Hamiltonian have a full-rank reduced density matrix associated with a sufficiently small subregion. In practice, the subregion should not exceed half the system. To complement the physical argument given above, we shall give an example in a frustrated quantum magnet in the next section.
@@ IV Icosahedral molecule
$H=J\sum_{{\rm bonds}\,\langle i,j\rangle}\sigma_{i}^{x}\sigma_{j}^{x}-h\sum_{{\rm sites}\,i}\sigma_{i}^{z}.$ (6)
@@ IV.1 Full-rank hypothesis and zero temperature phase diagram
As a first step, we checked that the FRH holds for subregions of adjacent spins with 1, 2 and 3 sites by obtaining the $2^{12}$ eigenstates via exact numerical diagonalization, see Appendix B. We then considered the fate of entanglement with varying magnetic field at zero temperature, so that the evolution is parameterized by $s=h$, see Fig. 2b. At h=0h\!=\!0, the 2-spin reduced density matrix is a full-rank separable state, which suggests the existence of a sudden-death value $h_{2}^{*}>0$ of the field below which the state is separable, and hence $\mathcal{E}(h<h_{2}^{*})=0$. We find $h_{2}^{*}\approx 0.6$. In contrast, the 3-spin reduced density matrix is not separable at small $h$ and we have $\mathcal{D}>0$ as shown in Fig. 2b. Turning to the GME between the spins on a triangular face, we find that it becomes very small for $h<0.65$, on the order of $10^{-6}$. Using the GTC, we are able to certify that the state is biseparable for $h\leqslant 0.25$. Although the certification becomes more difficult for larger field values, we believe that the true sudden death occurs near $0.65$.
@@ IV.4 Separation
with periodic boundary conditions such that $N+1\equiv N$. For our example, we randomly pick in the range [1,1][-1,1] the 3 exchange couplings $J_{i}$, and 3 components of the Zeemann magnetic field $h_{i}$: $\vec{J}=(-0.443,0.0938,0.915)$ and $\vec{h}=(0.930,-0.685,0.941)$. We work with $N=14$ sites. The groundstate is unique with energy gap $\Delta E=2.0007$. We have verified that the FRH holds for all $2^{14}$ eigenstates by computing the smallest eigenvalue of the reduced density matrix for a subregion made of 4 adjacent spins, as shown in Appendix B, which confirms the generic nature of $H_{\rm xyz}$.
@@ V.2 Quench
It would be desirable to study quenches in larger systems, where recent tensor network methods designed to track the reduced density matrix of small subregions [45] could, at least partially, allow one to study the fate of entanglement in a wider class of physical systems.
@@ VII Outlook
These new insights regarding the structure of entanglement in many-body states will have strong impact for quantum simulation and computation. For instance, efficient algorithms used to simulate quantum matter should incorporate the fact that distant particles are not entangled in spin/boson systems, which strongly constrains the variational space of many-body wave functions. Our work also paves the way for numerous outstanding research avenues. For instance, it will be essential to investigate how multipartite entanglement evolves in a plethora of realistic model Hamiltonians, and to determine the criteria that govern its demise under evolution. It would also be desirable to investigate systems which could escape the fate of entanglement. In parallel, we need a better understanding of violations of the FRH by certain systems with topological order or quantum many-body scars, for instance. Finally, we have also seen examples of how quantum matter out-of-equilibrium generates stronger multipartite entanglement than in equilibrium. This is the tip of iceberg: we expect that non-equilibrium dynamics can lead to rich entanglement structures that await to be discovered.